OpenStudy (anonymous):

how to solve 3x-5y+-20

OpenStudy (anonymous):

x= 6 2/3 y=4

OpenStudy (tester97):

subtract 3x from both sides -5y=-3x-20 divide both sides by -5 y=3/5x+4

OpenStudy (anonymous):

mybad -6 2/3

OpenStudy (tester97):

wait did i do that wrong?

OpenStudy (anonymous):

thanks guruslimm